Status Report

NASA Space Station On-Orbit Status 25 Feb 2004

By SpaceRef Editor
February 25, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 25 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

Preparations continue for the Orlan spacewalk tomorrow.  At 8:00am EST, the crew conducted a one-hour review of the updated EVA timeline, supported by tagup and discussion with ground specialists via S-band.  EVA-9 will begin with DC-1 EV hatch open at 4:20pm EST, for a duration of ~5h 35min.  While outside the station, the crew will rest during three eclipses (periods in Earth’s shadow) of 36 min. length each.

Other onboard preparations for the crew were as follows:

  • CDR Michael Foale configured the ECLSS (environment control & life support system) and TCS (thermal control system) in the U.S. segment (USOS) for the unmanned period.   [This essentially consisted in installing jumpers (fluid hoses) switching the TCS from MTL (moderate temperature loop) to LTL (low temperature loop) mode, in order to provide redundancy for critical avionics, thus minimizing the impacts of a depressurization in the USOS.  It also included commensurate reconfiguring of the computer OpsLAN (operations local area network) for the possibility of a Soyuz relocation from the DC-1 to the FGB.]
  • Foale also configured the USOS EPS (electrical power system) for the unmanned period.
  • FE Alexander Kaleri relocated three sensors of the Russian “Pille-MKS” radiation dosimeter system, placing one on a Service Module (SM) tray, the others in his and Mike Foale’s Orlan spacesuit pocket (left hip).
  • Mike meanwhile set up the video camera for covering the EVA, then installed the UOP-to-DPC (utility outlet panel to display & control panel) power bypass cable at the Lab RWS (robotic workstation).   [This allows ground control of the SSRMS (Space Station Remote Manipulator System) for TV coverage of the spacewalk tomorrow.]
  • Sasha worked on the Russian “Wiener” power laptop, installing an application and other files to configure it and its associated BRI SmartSwitch router for the Matryoshka radiation experiment to be installed outside.
  • The CDR refilled the disposable in-suit drink bags (DIDBs) and installed them in the Orlans.
  • Later in the afternoon, Mike Foale will also reconfigure the SSC (station support computer) network and transfer the procedures to microdrive for later reference.
  • Disassembly of the FGB-to-Node ventilation air ducts in the PMA-1 (pressurized mating adapter #1) by Sasha Kaleri is scheduled for 6:10pm tonight after deactivation of their internal TsV1 central ventilator by the ground.
  • Mike will then close all hatches in the USOS, completing the isolation of the segment from the rest of the station with the closure of the hatch between PMA-1 and FGB/Zarya.
  • Kaleri attended to his regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.
  • The Elektron electrolysis apparatus remains deactivated.   [Moscow plans two more troubleshooting tests on the failed Elektron, both post-EVA, viz.: an electrical checkout of its cabling, and a nitrogen purge of the internal lines of the Liquid Unit (BZh).  If both are unsuccessful, the BZh will be replaced around 3/6 with a new unit (#7) from spares.  The current unit (#6) was installed only last month (1/12), replacing the failed (redesigned) BZh #5, which had been installed in November 2002.  The BZh, with a diameter of 30 cm, height of 1.1 m and mass of 90 kg, supposedly has a certified life of one year.]

    Kaleri also conducted the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO), while Foale completed the weekly status checkup of autonomous Increment 8 payloads in the Lab.

    Shortly before sleep time tonight (9:15pm), Sasha will break out the “Urolux” equipment, setting it up for the Russian biochemical urine test (PZE MO-9), a standard requirement before and after Orlan-suited activities.   [The MO-9 analysis uses the sophisticated in-vitro diagnostic apparatus “Urolux” developed originally for the Mir program.]

    The crew worked out on TVIS treadmill, CEVIS cycle ergometer, RED exerciser and VELO stationary bike with load trainer.

    At 2:31pm, station attitude control was transferred from the U.S. to the RS motion control system, and at 2:36pm SM thrusters executed the 14-min. burn, which changed attitude from solar-oriented XPOP (x-axis perpendicular to orbit plane) to earth-oriented LVLH XVV (local vertical local horizontal, x-axis in velocity vector) in support of the EVA.  Control authority returns to the USOS at 3:40pm.   [During the EVA tomorrow, SM thrusters will be disabled, with attitude maintained by the U.S. CMGs.]

    After yesterday’s initial attempt by the crew to close and seal the Progress vestibule hatches, a leak was observed.  The hatches were re-closed, and the leak check this time indicated airtight seals.   [During the time of the problem the SM thrusters were kept disabled for an additional orbit, but ISS momentum was sufficiently low that a CMG desaturation was not required.]

    Today’s CEO targets (excluding Asia due to the current sleep cycle and North American and European targets due to winter weather conditions), constrained in XPOP attitude by flight rule to having the Lab science window only available for ~1/4 of each orbit when not facing in flight (“ram”) direction, were Bamako, Mali (nadir pass over this remote, hard-to-see city on the east banks of the Niger River.  Looking for the focus point of the radial road network), River patterns, S Amazonia (requested was a mapping swath right of track in a broad cloud-free zone.  Large delta-like river patterns of diverging courses have been detected recently on both sides of the Amazon, generated by major tributaries.  These unexpected features are not even completely mapped. They hold a rich history of landscape evolution), and Plankton, Argentine coast (Dynamic event.  The largest area of active plankton blooms lies in the southwest Atlantic from Patagonia to the Falklands. Looking south as ISS crossed the coast for structure in the plankton mass.  Overlapping images starting with an image that includes the coastline were requested).

    CEO images can be viewed at the websites.

    See also the website “Space Station Challenge” at

    U.S. and Russian Segment Status (as of today, 1:00pm EST).

    Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

    • Elektron O2 generator is Off.  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
    • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.8; ppO2 (mmHg) — 159.6; ppCO2 (mmHg) — 3.7;
    • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 21.0.
    • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 25.0.
    • Node:  Pressure (mmHg) — 756.61; temperature (deg C) — 23.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
    • U.S. Lab:  Pressure (mmHg) — 758.05; temperature (deg C) — 25.3; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
    • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 758.25; temperature (deg C) — 26.0; shell heater temp (deg C) — 23.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
    • PMA-1:  Shell heater temp (deg C) — 24.7
    • PMA-2:  Shell heater temp (deg C) — 12.1

    (n/a = data not available)

    Electrical Power Systems (EPS):

    • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position, with 47 deg bias for drag reduction.
    • SM batteries:  All batteries (8) are in “Partial Charge” mode.  
    • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
    • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

    Command & Data Handling Systems:

    • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
    • GNC-1 MDM is prime; GNC-2 is Backup.
    • INT-2 is operating; INT-1 is Off.
    • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
    • LA-1, LA-2 and LA-3 MDMs are all operating.
    • PL-1 MDM is Off; PL-2 MDM is Operational.
    • APS-1 (automated payload switch #1) and APS-2 are both On.
    • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
    • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

    Propulsion System:

    • Total propellant load available: 4070 kg (8972 lb) as of 2/12/04  [SM(755) + FGB(2656) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

    Attitude Control Systems:

    • 3 CMGs on-line (CMG-1 failed).
    • State vector source — U.S. SIGI-1 (GPS)
    • Attitude source — U.S. SIGI-1 (GPS)
    • Angular rate source — RGA-1

    Flight Attitude:

    • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

    Communications & Tracking Systems:

    • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
    • All other Russian communications & tracking systems are nominal.
    • *S-band is operating nominally (on string 2).
    • Ku-band is operating nominally.
    • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
    • Video subsystem operating nominally.
    • HCOR (high-rate communications outage recorder) is operating nominally.


    • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
    • MBS: KA power on both strings. 
    • MT: latched and mated at WS4. 
    • POA: KA power on both strings.
    • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

    ISS Orbit  (as of this morning, 6:38am EST [= epoch]):

    • Mean altitude — 366.1 km
    • Apogee — 370.3km
    • Perigee — 361.9 km
    • Period — 91.87 min.
    • Inclination (to Equator) —  51.629 deg
    • Eccentricity — 0.0006169
    • Orbits per 24-hr. day — 15.67
    • Mean altitude loss last 24 hours — 110 m
    • Revolutions since FGB/Zarya launch (Nov. ’98)  — 30063

    For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

    SpaceRef staff editor.